jiffy:

Chin canned a number of quart jars and a number of pint jars of tomatoes from his garden. A pint is 16 ounces and a quart is 32 ounces. The number of pints he canned is represented by p, and the number of quarts he canned is represented by q. Which equation models the scenario if Chin canned a total of 1,280 ounces?

jhonyy9:

a pint is 16 ounces and the number of pints canned repres. by p a quart is 32 ounces and the nr. of quarts repres. by q so from these result q = 2p using these details you need writing an equation how much p so pints and how much q so quarts you need assuming to get a total of 1280 ounces for example to get 80 ounces you can writing 2q + p = 80 chec it 2*32 +16 = 80 so 64 +16 =80 so 80 = 80 what is true hope helped understandably